OpenStudy (anonymous):

Which set of parametric equations represents the graph of the following rectangular equation using t = 1 - x. y = x2 + 2

OpenStudy (anonymous):

\[ x= 1-t\\ y=(1-t)^2 +2=1+t^2 -2t +2= t^2 -2t +3 \]

OpenStudy (anonymous):

Did you understand what I did?

OpenStudy (anonymous):

Yes, I more wanted clarification for my answer. Thanks sooo much!